Use the References to access important values if needed for this question. The following information is given for water at 1(please bold or circle the answer that I need to input)
Answer #1

Ti = 23.5 oC
Tf = 129.4 oC
here
Cl = 4.184 J/g.oC

Heat required to convert liquid from 23.5 oC to 100.0 oC
Q1 = m*Cl*(Tf-Ti)
= 45.9 g * 4.184 J/g.oC *(100-23.5) oC
= 14691.4884 J

Lv = 2259.0 J/g

Heat required to convert liquid to gas at 100.0 oC
Q2 = m*Lv
= 45.9g *2259.0 J/g
= 103688.1 J

Cg = 2.01 J/g.oC

Heat required to convert vapour from 100.0 oC to 129.4 oC
Q3 = m*Cg*(Tf-Ti)
= 45.9 g * 2.01 J/g.oC *(129.4-100) oC
= 2712.4146 J

Total heat required = Q1 + Q2 + Q3
= 14691.4884 J + 103688.1 J + 2712.4146 J
= 121092 J
= 121.1 KJ
Answer: 121.1 KJ